# Scaling Laws for Black-box Adversarial Attacks

### v3

@@ 5.1 Scaling Laws are General Principles
(b) Across different algorithms.
Figure 7: Ablations on untargeted settings and other algorithms. We mainly follow the experiment configurations in Sec. 3.1. For untargeted attacks, we set $\epsilon=2/255$. For other algorithms, we leverage cat as the target class and plot the average ASR over all target models. (a) It demonstrates our linear scaling laws work under untargeted settings. (b) It supports our claim that linear trends occur on variance-reduction algorithms, while MI-FGSM exhibits inferior performance due to gradient conflicts.
Untargeted Attack Setting. Untargeted attacks are similar to targeted attacks except for a different training objective from Eq. (2), and all our algorithms are obviously compatible with the untargeted setting:
$\max_{\bm{x}}\frac{1}{T}\sum_{i=1}^{T}\mathcal{L}(f_{i}(\bm{x}),\bm{y}_{\text{nat}}),\text{s.t.}\left\lVert\bm{x}-\bm{x}_{\text{nat}}\right\rVert_{\infty}\leq\epsilon,$ (11)
Our experiment results in Fig. 7(a) serve as a clear sign that scaling laws remain valid over untargeted settings. We mainly follow our settings in Sec. 3.1, while we set $\epsilon=2/255$, and measure the results on 3 target models. Taking an extremely small $\epsilon$ is necessary for demonstrating our linear scaling laws, otherwise the attack success rate will easily reach $100\%$ and therefore come to a plateau. The results validate our claim that scaling laws exist regardless of specific attack targets.
Across Different Attack Algorithms. Besides SSA-CWA, we also provide experimental validations of different attack algorithms in Fig. 7(b). As stated in Sec. 2.2, severe gradient conflicts can be observed when we iteratively update the input image $\bm{x}$ if we directly perform backward propagation over the sum of output logits (e.g., as in MI-FGSM). Gradient conflicts prevent the input $\bm{x}$ from being updated properly and lead to missing information of model weaknesses. However, advanced algorithms, such as SVRE [92] and CWA [9] address this issue by leveraging optimizers that can reduce the variance of gradients. Our experiment results support our claim that scaling laws occur only on such algorithms. Note that the attack success rate of MI-FGSM also becomes higher with increasing ensemble cardinality, further validating that our theoretical insight from Theorem 2 is regardless of specific attack methods.
5.2 Visualizations of Adversarial Perturbations
Figure 8: Visualization of adversarial perturbations. A clear trend of the increasing semantic interpretability of adversarial perturbations over different target objects.
As shown in Fig. 8, the semantic meaning of our adversarial perturbations also increases as the cardinality of the model ensemble grows. This suggests that the attack success rates are closely related to the interpretability of the perturbations, aligning with human observations and preferences [15]. Since the ultimate goal of adversarial attacks is to generate high-quality adversarial examples, it is preferable to have interpretable perturbations, as this reinforces the idea that the scaling effectively helps learn common features rather than merely converging to local optima.
As people traditionally believe images lie on a low dimensional manifold of desirable properties [31], this interpretability of adversarial examples can be also understood within the feature space of images, where human-designed attack algorithms aim to eliminate the non-robust features of images and shift the image towards the robust features of the target class [45]. From this perspective, our approach of enlarging model ensembles seeks to provide more information about the manifold itself, offering more accurate representations of the image features.
When we enlarge $\epsilon$ from $8/255$ to $16/255$ or even larger, the effect of the natural image goes down to 0, and the adversarial attack task degenerates to an image generation task. We claim that our method of scaling model ensemble still performs well under the degenerated image generation task, where the visualized results are shown in Appendix F.
5.3 Fitting Error of Scaling Law Curve
To comprehensively demonstrate the universality of our scaling laws, we provide the MSE fitting error from Fig. 3 in Table.3. It can be observed that the small fitting error acts as a quantitative validation of our scaling laws. Thus, our scaling laws remain a scalable way to enhance the transferability of adversarial examples.
Targets Cat Car Ship Deer Bird
Metric ASR Loss ASR Loss ASR Loss ASR Loss ASR Loss
Fitting Error 0.0040 0.34 0.0040 0.17 0.0038 0.46 0.0025 0.05 0.0023 0.27
Table 3: The averaged MSE fitting error of our curves in Fig. 3. Note that ASR usually has range $[0,1]$, while loss has range $[0,10]$. The fitting error is averaged on all test models, not only the test models that are fitted with lines for demonstration.
@@ 6 Conclusion
In this work, we investigate the scaling laws for black-box adversarial attacks. Theoretically, we provide asymptotic analysis on the convergence rate of generation error and adversarial examples in Sec. 3.1. Empirically, we show that the attack success rate and the loss are converging exponentially. We verify this by comprehensive experiments over traditional image classifiers (Sec. 3.2), defended models like AT and DiffPure (Sec. 3.3), and modern MLLMs (Sec. 3.4 and Sec. 4) using various optimizers (Sec. 5). These experimental results demonstrate the universality of scaling laws in transfer attacks.
Leveraging our insights into scaling laws, we emphasize further research on evaluating black-box weaknesses of MLLMs by focusing on large-scale experiments, rather than focusing on designing intricate algorithms on a small set of surrogates, which may deepen the understanding of the weaknesses of large models, ultimately contributing to the development of more robust and trustworthy multimodal large foundation models.
@@
[…]
@@ Appendix B Experiment Settings for Scaling Laws
In this section we list our models that are included in Sec. 3.2. We select these models based on their diversity in architectures, training methods, and training datasets. Results are shown in Table 4 and Table 5, where all of our models come from open-source platforms such as Torchvision [66] and Timm [84]. Note that default in the table means normally trained on ImageNet-1K dataset.
@@ Appendix C SSA-CWA Algorithm Declaration
We provide the pseudo-code version of SSA-CWA implementation in Algorithm 1. Their combination is induced from  [63] and  [9], while implementations of other algorithms (e.g., MI-FGSM, MI-CWA) can be found in their original papers.
@@
5:     for i = 1 todo
6:         for j = 1 todo $\triangleright$ Spectrum transformation
@@
Appendix D Test Prompts as Metrics
As VLMs and MLLMs output prompts rather than logits, we evaluate the attack efficacy by mainly focusing on the “classification” results of the test models, which is leveraging a GPT4 assistant to judge the content of the test models’ descriptions. We develop Metric 1 and Metric 2, focusing on different aspects of the descriptions.
Fig. 9(a) and Fig. 9(b) display our prompts for Metric 1 and Metric 2 respectively. Metric 1 follows the settings in previous works [98], while Metric 2 is our customized metric to capture image features, as stated in Sec. 4.2. By comparing the results in Fig. 6 and Table. 2, Metric 2 have a 10% higher success rate than Metric 1 under $\epsilon=8/255$ cases. This phenomenon is natural since some images of high noise will be still considered as success in Metric 2, which is more aligned with our practical approaches [26].
(a) Prompts for Metric 1.
(b) Prompt for Metric 2.
Figure 9: Prompts of two metrics. (a) Besides prompt to test the adversarial example contain a target label, we also use a prompt to judge the image quality, ensuring our successful attacks are all of small noise and highly transferable. (b) The prompt concentrates on semantic meanings of the image descriptions. To consider the attack successful, it only needs the target to appear in the description.
@@ Appendix E Experiment Settings for Attacking MLLMs
Figure 10: Visualizations of adversarial examples under $\epsilon=64/255$. The adversarial examples capture basic information of the target class, such as the fur of cats, tires of vehicles, cargo on ships, antlers of deer, and feathers of birds.
@@
In Table. 6 we list the 12 CLIPs that are collected for our model ensemble, which are all open-source models selected from OpenAI [70], OpenCLIP [12], and Timm [84].
Appendix F Degenerated Image Generation Task
As stated in Sec. 5.2, when $\epsilon$ goes larger, the adversarial attack task degenerates to an image generation task. Previous works [9, 1, 82] do not ever think about this field, but we believe that adversarial attack task and image generation task are indeed homogenous. We display our adversarial examples generated from 64 image classifiers (stated in Appendix B) under $\epsilon=64/255$ in Fig. 10. From the visualizations we observe that our methods can capture the key features of the target class and reproduce them based on the original textures of images. This visualization further supports our methods are really attacking common weakness instead of running into local optima. We leave further explorations and analyses as future work.
Appendix G Related Work
G.1 Black-Box Attacks
@@
Deep learning models are widely recognized for their vulnerability to adversarial attacks [73, 35]. Extensive research has documented this susceptibility, particularly in image classification contexts [2, 7], under two principal settings: white-box and black-box attacks. In the white-box setting, the adversary has complete access to the target model [65]. Conversely, in the black-box setting, such information is obscured from the attacker. With limited access to the target model, black-box attacks typically either construct queries and rely on feedback from the model [25, 44] or leverage the transferability of predefined surrogate models [60]. Query-based attacks often suffer from high time complexity and substantial computational costs, making transfer-based methods more practical and cost-effective for generating adversarial examples, as they do not necessitate direct interaction with target models during the training phase.
@@
Within the realm of transfer-based attacks, existing methods can generally be categorized into three types: input transformation, gradient-based optimization, and ensemble attacks. Input transformation methods employ data augmentation techniques before feeding inputs into models, such as resizing and padding [89], translations [23], or transforming them into frequency domain [63]. Gradient-based methods concentrate on designing optimization algorithms to improve transferability, incorporating strategies such as momentum optimization [22, 57], scheduled step sizes [33], and gradient variance reduction [81].
G.2 Ensemble Attacks
@@
Inspired by model ensemble techniques in machine learning [87, 19, 34], researchers have developed sophisticated adversarial attack strategies utilizing sets of surrogate models. In forming an ensemble paradigm, practitioners often average over losses [22], predictions [60], or logits [22]. Additionally, advanced ensemble algorithms have been proposed to enhance adversarial transferability [55, 9, 92, 8]. However, these approaches tend to treat model ensemble as a mere technique for improving transferability without delving into the underlying principles.
@@
Prior studies [41] indicate that increasing the number of classifiers in adversarial attacks can effectively reduce the upper bound of generalization error in Empirical Risk Minimization. Recent research [94] has theoretically defined transferability error and sought to minimize it through model ensemble. Nonetheless, while an upper bound is given, they do not really scale up in practice and have to train neural networks from scratch to meet the assumption of sufficient diversity among ensemble models [59].
G.3 Scaling Laws
@@
In recent years, the concept of scaling laws has gained significant attention. These laws describe the relationship between model size – measured in terms of parameters, data, or training FLOPs – and generalization ability, typically reflected in evaluation loss [48]. Subsequent studies have expanded on scaling laws from various perspectives, including image classification [3, 38], language modeling [71] and mixture of experts (MoE) models [30], highlighting the fundamental nature of this empirical relationship.
@@
In the domain of adversarial attack research, recent work has attempted to derive scaling laws for adversarially trained defense models on CIFAR-10 [50, 5]. However, the broader concept of scaling remains underexplored. While some preliminary observations on the effect of enlarging model ensembles can be found in prior studies [94], these effects have not been systematically analyzed. Recent work Wei et al. [83] also explore the scaling law of attack success rate when jailbreaking large language models [99, 10] by increasing number of in-context demonstrations In this paper, however, we focus on the empirical scaling laws within model ensembles, and verify it through extensive experiments on open-source pretrained models. Motivated by this perspective, we unveil vulnerabilities within commercial LLMs by targeting their vision encoders.

### v4

@@ 5.1 Ablations on the Scaling Laws’ Generality
(b) Across different optimizers.
Figure 6: Ablations on the generality of the log-linear scaling laws. (a) The laws persist in the untargeted setting (Eq. 3). (b) The laws hold for advanced optimizers (MI-CWA, VMI-CWA, MI-SVRE) that mitigate gradient conflict. In contrast, the naïve logit ensemble strategy fails to scale, supporting our hypothesis in Sec. $2.$
5.1.1 Untargeted Attack Setting.
We first validate the scaling laws in the untargeted setting (defined by Eq. 3). To do this, we follow the core experimental settings from Sec. 3.2, using the SSA-CWA algorithm on the same 64-model surrogate pool and NIPS17 dataset. We measure the average ASR against 3 representative target models from Family 1 (Swin-B, ConvNeXt-B, and ResNeXt50). As the untargeted setting is a significantly simpler task, We use a small $\epsilon=2/255$ to avoid the ASR quickly saturating at 100%, which would obscure the underlying trend. As shown in Fig. 6(a), even under untargeted conditions, a clear log-linear trend persists. This confirms that the scaling laws are a general principle of ensemble attacks, independent of the specific attack objective.
5.1.2 Across Different Attack Algorithms.
We now provide a tailored experiment to validate our hypothesis from Sec. 2.2: that the log-linear scaling laws are unlocked only by advanced ensemble optimizers (like CWA or SVRE) that resolve gradient conflict, while naïve ensemble strategies fail to scale.
Experimental Design. To test this, we design two families of attack algorithms and evaluate their scaling performance. All experiments follow the protocol in Sec. 3.2 (targeting sea lion) and results are averaged over our standard target classifiers (Family 1 of Sec. 3.2).
  • Family 1 (Logit Ensemble Strategies): We test the logit averaging strategy, which we hypothesize will fail to scale. We combine the strategy with various powerful gradient optimizers and input transformations: MI-FGSM [dong2018boosting], DI-MI-FGSM [xie2019improving], and VMI-FGSM [wang2021enhancing].
  • Family 2 (Advanced Ensemble Strategies): We then test advanced gradient optimizers. This family includes MI-CWA [chen2023rethinking], VMI-CWA [wang2021enhancing, chen2023rethinking], and MI-SVRE [xiong2022stochastic].
Results and Analysis. The results are shown in Fig. 6(b). All algorithms in Family 1, which use logit averaging, fail to exhibit the log-linear scaling law. While performance can be boosted by components like DI or VMI, their ASR stagnates at a low value. Conversely, algorithms in Family 2 demonstrate a clear and persistent log-linear scaling law. This behavior mirrors the results of our main SSA-CWA algorithm (from Sec. 3), confirming that the scaling phenomenon is robust across different advanced optimizers, so long as they resolve the conflict. This experiment confirms our claim: the ability to scale an ensemble attack is contingent on the ensemble optimization strategy.
5.2 The Semantic Nature of Scaled Perturbations
Figure 7: Visualization of adversarial perturbations as $T$ scales. We take 5 labels from CIFAR10 [krizhevsky2009learning] to create adversarial examples as they are relatively easy for human eyes to recognize. As the ensemble cardinality grows (top to bottom), the perturbations evolve from unstructured noise ($T=1$) into semantically meaningful patterns that resemble features of target classes ($T=64$).
Beyond attack success rates, we investigate what the ensemble attack learns as $T$ increases. As shown in Fig. 7, scaling the ensemble fundamentally changes the nature of the resulting perturbation. At $T=1$, the perturbation is noisy and model-specific. However, as $T$ increases to 64, the perturbation evolves to contain clear, interpretable, and semantic features of the target class (e.g., textures for ‘Cat’, an owl-like shape for ‘Bird’).
This is a critical finding that provides a qualitative explanation for the scaling laws. First, scaling $T$ forces the optimizer to discard model-specific, ‘non-robust features’ [ilyas2019adversarialexamplesbugsfeatures] and instead find a ‘common weakness’ solution that fools all models. Next, our visualization strongly suggests this common weakness is not just a mathematical artifact but is tied to the semantic essence of the target class. The attack learns to inject robust features of the target [dai2022fundamental].
This reinforces that our scaling approach is a principled method to distill the shared, transferable features of a concept from a diverse set of models. As noted by [fawzi2018adversarial], images lie on a low-dimensional manifold; our scaling approach appears to provide a more accurate and robust representation of this manifold for the target class.
@@ 6 Conclusion
In this work, we investigate scaling laws for black-box adversarial attacks. We find that while gradient conflict stalls naïve ensembling, advanced optimizers (e.g., CWA) unlock a robust log-linear scaling law: the Attack Success Rate increases linearly with $\log T$. We empirically verify this law’s universality across standard classifiers (Sec. 3.3), SOTA defenses , and MLLMs (Sec. 3.4), and show that scaling distills semantic features (Sec. 5.2). Applying this methodology to SOTA MLLMs (Sec. 4) serves as a powerful benchmark. The scaled attack achieves 85.0% ASR on GPT-4o, while Claude-3.5-Sonnet shows exceptional resilience at only 31.0% ASR. Our findings urge a shift in focus for robustness evaluation: from designing intricate algorithms on small ensembles to understanding the principled and powerful threat of scaling, which is crucial for building and verifying genuinely robust foundation models.
@@
$\sqrt{T}(\hat{\bm{x}}-\bm{x}^{*})\overset{d}{\to}(\nabla^{2}\mathbb{L}(\bm{x}^{*}))^{-1}\mathcal{N}(0,\mathrm{Cov}(\mathbb{L}(f_{i}(\bm{x}^{*}))))$
$\overset{d}{=}\mathcal{N}(0,\nabla^{2}\mathbb{L}(\bm{x}^{*})^{-1}\mathrm{Cov}(\mathbb{L}(f_{i}(\bm{x}^{*}))\nabla^{2}\mathbb{L}(\bm{x}^{*})^{-1}).$
@@ Appendix B Experiment Settings for Scaling Laws
In this section we list our models that are included in Sec. 3.2. We select these models based on their diversity in architectures, training methods, and training datasets. Results are shown in Table 2 and Table 3, where all of our models come from open-source platforms such as Torchvision [torchvision2016] and Timm [rw2019timm]. Note that default in the table means normally trained on ImageNet-1K dataset.
@@ Appendix C SSA-CWA Algorithm Declaration
We provide the pseudo-code version of SSA-CWA implementation in Algorithm 1. Their combination is induced from  [long2022frequency] and  [chen2023rethinking], while implementations of other algorithms (e.g., MI-FGSM, MI-CWA) can be found in their corresponding papers.
@@
5:  for i = 1 to N do
6:   for j = 1 to M do $\triangleright$ Spectrum transformation
@@
Appendix D Evaluation Metrics Details
As MLLMs output generative text rather than classification logits, assessing attack success requires parsing natural language descriptions. In this work, we employ two distinct evaluation paradigms tailored to the specific experimental goals in Sec. 3 and Sec. 4.
In Sec. 3.4.2 (Family 3: MLLMs), we evaluate the scaling laws using a string-matching approach. Since MLLMs often produce descriptive text rather than the precise ImageNet class names, strict exact matching can lead to false negatives. Therefore, we adopt a relaxed keyword map. An attack is considered successful if the MLLM’s output contains any of the keywords associated with the target label.
Table 4 lists the target labels used in our experiments and their corresponding relaxed keywords.
Target Label (ImageNet) Accepted Keywords (Case-Insensitive)
Sea Lion sea lion, seal, marine mammal
Barometer barometer, gauge, measuring instrument, clock
Swing swing, playground, seat
Pickelhaube pickelhaube, helmet, spiked helmet, military hat
Fig fig, fruit, plant
Lock lock, padlock, dial
Computer computer, pc, laptop, desktop, phone, device
Snake snake, serpent, reptile, python, cobra
Table 4: The relaxed keyword map used for evaluating MLLMs in Sec. 3. We map the specific ImageNet class label to a set of semantically equivalent or broader keywords.
In Sec. 4, to rigorously benchmark SOTA MLLMs, we adopt the LLM-as-Judge framework inspired by MultiTrust [zhang2024benchmarking]. We use GPT-4o as the judge to evaluate the quality and success of the attack.
The evaluation is strict: an attack is successful only if the description mentions the target and does not mention artifacts of the attack (noise, distortion). Fig. 8 presents the exact system prompt used in our experiments.
Prompt for GPT-4o Judge Correctness: Given a chatbot’s description of a scene: [response]. If it is stated in the description that there is a/an [label]? Please answer ’Yes.’ or ’No.’. Prompt for GPT-4o Judge Image Quality: Given a chatbot’s description of an image: [response]. Does the description mention factors that affect the judgment of the image content, such as low image quality, noise, or modification? Please answer ’Yes.’ or ’No.’.
Figure 8: The GPT-4o system prompt used for the unified evaluation metric in Sec. 4. This metric ensures that successful attacks are both effective (semantics changed) and stealthy (no obvious noise detected by the model).
@@ Appendix E Experiment Settings for Attacking MLLMs
In Table. 5 we list the 12 CLIPs that are collected for our model ensemble, which are all open-source models selected from OpenAI [radford2021learningtransferablevisualmodels], OpenCLIP [cherti2023reproducible], and Timm [rw2019timm].
Appendix F Related Work
F.1 Black-Box Attacks
@@
Deep learning models are widely recognized for their vulnerability to adversarial attacks [szegedy2013intriguing, goodfellow2014explaining]. Extensive research has documented this susceptibility, particularly in image classification contexts [athalye2018obfuscated, carlini2017towards], under two principal settings: white-box and black-box attacks. In the white-box setting, the adversary has complete access to the target model [madry2017towards]. Conversely, in the black-box setting, such information is obscured from the attacker. With limited access to the target model, black-box attacks typically either construct queries and rely on feedback from the model [dong2021query, ilyas2018black] or leverage the transferability of predefined surrogate models [liu2016delving]. Query-based attacks often suffer from high time complexity and substantial computational costs, making transfer-based methods more practical and cost-effective for generating adversarial examples, as they do not necessitate direct interaction with target models during the training phase.
@@
Within the realm of transfer-based attacks, existing methods can generally be categorized into three types: input transformation, gradient-based optimization, and ensemble attacks. Input transformation methods employ data augmentation techniques before feeding inputs into models, such as resizing and padding [xie2019improving], translations [dong2019evading], or transforming them into frequency domain [long2022frequency]. Gradient-based methods concentrate on designing optimization algorithms to improve transferability, incorporating strategies such as momentum optimization [dong2018boosting, lin2019nesterov], scheduled step sizes [gao2020patch], and gradient variance reduction [wang2021enhancing].
F.2 Ensemble Attacks
@@
Inspired by model ensemble techniques in machine learning [wortsman2022model, dietterich2000ensemble, gontijo2021no], researchers have developed sophisticated adversarial attack strategies utilizing sets of surrogate models. In forming an ensemble paradigm, practitioners often average over losses [dong2018boosting], predictions [liu2016delving], or logits [dong2018boosting]. Additionally, advanced ensemble algorithms have been proposed to enhance adversarial transferability [li2023making, chen2023rethinking, xiong2022stochastic, chen2023adaptive]. However, these approaches tend to treat model ensemble as a mere technique for improving transferability without delving into the underlying principles.
@@
Prior studies [huang2023t] indicate that increasing the number of classifiers in adversarial attacks can effectively reduce the upper bound of generalization error in Empirical Risk Minimization. Recent research [yao2025understandingmodelensembletransferable] has theoretically defined transferability error and sought to minimize it through model ensemble. Nonetheless, while an upper bound is given, they do not really scale up in practice and have to train neural networks from scratch to meet the assumption of sufficient diversity among ensemble models [liu2019deep].
F.3 Scaling Laws
@@
In recent years, the concept of scaling laws has gained significant attention. These laws describe the relationship between model size – measured in terms of parameters, data, or training FLOPs – and generalization ability, typically reflected in evaluation loss [kaplan2020scalinglawsneurallanguage]. Subsequent studies have expanded on scaling laws from various perspectives, including image classification [bahri2024explaining, hestness2017deep], language modeling [sharma2022scaling] and mixture of experts (MoE) models [sun2024hunyuanlargeopensourcemoemodel], highlighting the fundamental nature of this empirical relationship.
@@
In the domain of adversarial attack research, recent work has attempted to derive scaling laws for adversarially trained defense models on CIFAR-10 [krizhevsky2009learning, bartoldson2024adversarialrobustnesslimitsscalinglaw]. However, the broader concept of scaling remains underexplored. While some preliminary observations on the effect of enlarging model ensembles can be found in prior studies [yao2025understandingmodelensembletransferable], these effects have not been systematically analyzed. Recent work wei2023jailbreak also explore the scaling law of attack success rate when jailbreaking large language models [zou2023universal, chen2025towards] by increasing number of in-context demonstrations In this paper, however, we focus on the empirical scaling laws within model ensembles, and verify it through extensive experiments on open-source pretrained models. Motivated by this perspective, we unveil vulnerabilities within commercial LLMs by targeting their vision encoders.